# MMIDR: Teaching Large Language Model to Interpret Multimodal Misinformation via Knowledge Distillation

### v1

@@ Abstract.
Automatic detection of multimodal misinformation has gained a widespread attention recently. However, the potential of powerful Large Language Models (LLMs) for multimodal misinformation detection remains underexplored. Besides, how to teach LLMs to interpret multimodal misinformation in cost-effective and accessible way is still an open question. To address that, we propose MMIDR, a framework designed to teach LLMs in providing fluent and high-quality textual explanations for their decision-making process of multimodal misinformation. To convert multimodal misinformation into an appropriate instruction-following format, we present a data augmentation perspective and pipeline. This pipeline consists of a visual information processing module and an evidence retrieval module. Subsequently, we prompt the proprietary LLMs with processed contents to extract rationales for interpreting the authenticity of multimodal misinformation. Furthermore, we design an efficient knowledge distillation approach to distill the capability of proprietary LLMs in explaining multimodal misinformation into open-source LLMs. To explore several research questions regarding the performance of LLMs in multimodal misinformation detection tasks, we construct an instruction-following multimodal misinformation dataset and conduct comprehensive experiments. The experimental findings reveal that our MMIDR exhibits sufficient detection performance and possesses the capacity to provide compelling rationales to support its assessments. [^1] Code and data are available at https://github.com/wishever/MMIDR
@@ 3.1.2. Evidence Retrieval
[…]
@@ 3.2. Rationales Elicitation
Now give you a multimodal post and the corresponding textual evidences and image evidences, these evidences may be relevant or irrelevant. Each post is a tuple of (text, image_ocr, image_caption). Evidences are a list of sentences, which are split by ”¡and¿”. ## Three-way classification scheme Each post is labeled in three-way classification scheme: [0: ”non-rumor”, 1: ”rumor”, 2: ”unverified”]. ## Fine-grained labels ¡Non-rumor — Rumor — Unverified fine-grained labels¿ ## Post Post:[ text: {post_text}, image ocr: {image_ocr}, image caption: {image_caption}] ## Evidences Textual_evidence: {textual_evidence} Image_evidence: {visual_evidence} ## Label label: ¡Non-rumor — Rumor — Unverified¿ ## Explanation Now, you are required to explain why the post is labeled as ¡Non-rumor — Rumor — Unverified¿ according to the evidences, and decide which fine-grained labels are suitable for the post.
@@ 4.1.1. Datasets
[…]
@@ 4.2. RQ1: Zero-shot Experiments
In this section, we perform zero-shot experiments to verify whether large language models inherently possess enough background knowledge to detect multimodal misinformation. Specifically, we evaluate both the teacher model ChatGPT (gpt-3.5-turbo) (Ouyang et al., 2022) and the student model MiniGPT-v2 (Chen et al., 2023) on the test set of $\text{MR}_{llm}^{2}$. The results are presented in Table 3.
@@ 4.3. RQ2: Performance Comparison and Ablation Study
Input Post
@@
天津一女子在路上扶起了摔倒的老人帮忙送去救治并垫付3000元,老人清醒后指认是女子撞到了他,老人的子女要求女子再拿10000元。女子报了警… (A woman in Tianjin helped an elderly person who had fallen on the road, took him for medical treatment, and paid 3,000 yuan. After the elderly person regained consciousness, he accused the woman of causing the accident. The elderly person’s children demanded an additional 10,000 yuan from the woman. The woman reported the incident to the police.)
While India is in denial about border tensions in Ladakh, China linked accounts shared this photo with the caption India China border, in which Indian soldiers are seen lying on the ground with hands and legs tied up with rope.
Instruction
@@
Now give you a multimodal post and the corresponding textual evidences… image ocr: 老人跌倒路边一直求救女子暖心扶起帮老人找家我拉你一下好不好@历史课课代表, (The elderly person fell by the roadside, continuously calling for help. The compassionate woman helped the elderly person up and assisted in finding his way home, saying, ”Let me give you a hand, is that okay?” @HistoryClassRep) image caption:an old man lying on the ground with his head down… Evidences… ¡and¿ 雷音振长空01-0220:57 微博轻享版任何值得称道的正义之光,都应建立在事实之上。可惜的是,这个被这么多大v转发的新闻是假的 (Any commendable light of justice should be built upon facts. Unfortunately, the news widely shared by many influencers is fake.) … Now,you are required to explain…
Now give you a multimodal post and the corresponding textual evidences… image ocr: @Lz1277 87829, image caption: soldiers lay on the ground while others stand around them. Evidences… ¡and¿ 72 Indian soldiers injured, captured in clash with Chinese Army… ¡and¿ Footage claims to show ’Indian troops beating a captured Chinese soldier’ amid border tension… ## Explanation: Now, you are required to explain and label the post in three-way classification scheme: [”non-rumor”, ”rumor”, ”unverified”].
gpt-3.5-turbo
The post is labeled as rumor because the textual evidences show conflicting information and the image caption does not directly support the text. The textual evidence includes multiple instances of the post being shared on social media, but also includes a comment suggesting that the news is fake. Additionally, the image OCR does not provide any relevant information to confirm the story in the text. Fine-grained labels that are suitable for the post include ”Misleading and False” and ”Fake News” as the story is being disputed and there is a suggestion that it is not true. Therefore, the post is labeled as ¡label¿ rumor ¡/label¿.
@@
The post is labeled as a rumor because the textual evidence mentions that the photo was circulated among Chinese military websites, and there are conflicting reports about the incident, with some claiming that Indian soldiers were captured and others claiming that Indian troops beat a captured Chinese soldier. This indicates that there is uncertainty and conflicting information surrounding the photo, leading to its classification as a rumor. As for the fine-grained labels, ”Altered Image” and ”Misleading” would be suitable for the post, as the photo is being shared with a misleading caption and there are conflicting reports about the incident depicted in the image. Therefore, the post is labeled as ¡label¿ rumor ¡/label¿.
$\text{MMIDR}_{\text{LLaMA2}}$
@@
The post is labeled as rumor because the textual and image evidences do not provide clear and credible information about the incident. The textual evidence includes repeated mentions of ”Sina Visitor System,” which is not relevant to the content of the post. The image evidence includes a screenshot of a text message in Chinese, which does not provide any relevant information about the incident. Additionally, the image caption describes an old man lying on the ground, which is not directly related to the claim made in the post. Fine-grained labels that are suitable for the post include ”Misleading,” ”Lacks Context,” and ”Unverified,” as the post lacks clear and credible information, and the image evidence does not support the claim made in the text. Therefore, the post is labeled as ¡label¿ rumor ¡/label¿.
The post is labeled as a rumor because the textual evidence suggests that the photo of Indian soldiers being tied up with rope is being circulated among Chinese military websites, but there is no concrete evidence to support this claim. Additionally, the image OCR and caption do not provide any relevant information to verify the claim made in the text. Fine-grained labels that are suitable for the post include: Altered Image, Misleading, Misattributed, and False. Therefore, the post is labeled as ¡label¿ rumor ¡/label¿.
$\text{MMIDR}_{\text{MiniGPT-v2}}$
The post is labeled as a rumor because the textual evidence includes multiple screenshots of Chinese social media posts and text messages, but there is no concrete evidence or credible source to support the claim that the woman in the post helped an old man and paid 3000 yuan for his medical expenses. The image caption of an old man lying on the ground with his head down also does not provide any concrete evidence to support the claim. Fine-grained labels that are suitable for the post include ”Misleading”, ”Lacks Context”, and ”Unverified”. These labels reflect the lack of concrete evidence and the potential for misleading information in the post. Therefore, the post is labeled as ¡label¿ rumor ¡/label¿.
@@
The post is labeled as rumor because the textual and image evidences suggest that there is a claim of Indian soldiers being tied up and beaten by Chinese soldiers, which is not verified. The textual evidence mentions a photo circulating among Chinese military websites showing PLA troops bringing down several Indian soldiers at a close-quarters fight at Pangong Lake border area with India. The image evidence does not directly support the claim, but it does not refute it either. Fine-grained labels that are suitable for the post include ”Misleading”, ”Misrepresented”, and ”Unverified”. These labels reflect the lack of verification and the potential for misleading information in the post. Therefore, the post is labeled as ¡label¿ rumor ¡/label¿
@@ 4.4. RQ3: Faithful Detection and Interpretable Results
To evaluate the ability of LLMs in producing high-quality and coherent text explanations for multimodal misinformation, we conducted a case study, as detailed in Table 5. We specifically chose one Chinese example and one English example for in-depth analysis.

### v2

@@ Abstract.
Automatic detection of multimodal misinformation has gained a widespread attention recently. However, the potential of powerful Large Language Models (LLMs) for multimodal misinformation detection remains underexplored. Besides, how to teach LLMs to interpret multimodal misinformation in cost-effective and accessible way is still an open question. To address that, we propose MMIDR, a novel framework for multimodal misinformation interpretation and distillation reasoning, aiming to teach LLMs in providing fluent and high-quality textual explanations for their decision-making process. To convert multimodal misinformation into an appropriate instruction-following format, we present a data augmentation perspective and pipeline. This pipeline consists of a visual information processing module and an evidence retrieval module. Subsequently, we prompt the proprietary LLMs with processed contents to extract rationales for interpreting the authenticity of multimodal misinformation. Furthermore, we design an efficient knowledge distillation approach to distill the capability of proprietary LLMs in explaining multimodal misinformation into open-source LLMs. To explore several research questions regarding the performance of LLMs in multimodal misinformation detection tasks, we construct an instruction-following multimodal misinformation dataset and conduct comprehensive experiments. The experimental findings reveal that our MMIDR exhibits sufficient detection performance and possesses the capacity to provide compelling rationales to support its assessments.
@@ 3.1.2. Evidence Retrieval
Similarly, we employ the textual content $T$ from the input post as a query and leverage the Google Programmable Search Engine [^3] https://programmablesearchengine.google.com to obtain visual evidence denoted as $E_{V}$. The search engine generates a collection of images corresponding to the query text. For these visual evidentiary elements, we implement the aforementioned visual information processing techniques (§$\S$3.1.1), encompassing OCR extraction and image caption generation. Hence, a piece of visual evidence $E_{V}=\{Image\_Title,Image\_OCR,Image\_Caption\}$ represents the corresponding image’s title, OCR extraction results, and image caption generation results, respectively.
@@ 3.2. Rationales Elicitation
Labeling Prompt Template Now give you a multimodal post and the corresponding textual evidences and image evidences, these evidences may be relevant or irrelevant. Each post is a tuple of (text, image_ocr, image_caption). Evidences are a list of sentences, which are split by ”¡and¿”. ## Three-way classification scheme Each post is labeled in three-way classification scheme: [0: ”non-rumor”, 1: ”rumor”, 2: ”unverified”]. ## Fine-grained labels ¡Non-rumor — Rumor — Unverified fine-grained labels¿ ## Post Post:[ text: {post_text}, image ocr: {image_ocr}, image caption: {image_caption}] ## Evidences Textual_evidence: {textual_evidence} Image_evidence: {visual_evidence} ## Label label: ¡Non-rumor — Rumor — Unverified¿ ## Explanation Now, you are required to explain why the post is labeled as ¡Non-rumor — Rumor — Unverified¿ according to the evidences, and decide which fine-grained labels are suitable for the post.
@@ 4.1.1. Datasets
We employ ChatGPT (gpt-3.5-turbo) (Ouyang et al., 2022) as the teacher model to elucidate the rationales behind labeling multimodal misinformation based on evidence and determining fine-grained labels applicable to each post. The result instruction-following dataset is denoted as $\text{MR}_{llm}^{2}$. The statistical features and labels distribution of $\text{MR}_{llm}^{2}$ are presented in Table 2. Additionally, we graphically depict the length of instances and generated output with varying quantities of textual evidence ($m$) and visual evidence ($n$) obtained through the evidence retrieval module (§$\S$3.1.2), as illustrated in Figure 2. Specifically, we instantiate instances with different quantities of evidence using the instruction template, concatenating the rationales generated by the teacher model for illustration. We provide visualizations of the lengths of all instances, which serve as a basis for subsequent experiments and as references in hyperparameter settings.
@@ 4.2. RQ1: Zero-shot Experiments
In this section, we perform zero-shot experiments to verify whether large language models inherently possess enough background knowledge to detect multimodal misinformation. Specifically, we evaluate both ChatGPT (gpt-3.5-turbo) (Ouyang et al., 2022) and MiniGPT-v2 (Chen et al., 2023) on the test set of $\text{MR}_{llm}^{2}$. The results are presented in Table 3.
@@ 4.3. RQ2: Performance Comparison and Ablation Study
Input Post [Uncaptioned image] [Uncaptioned image]
@@
{CJK}UTF8gbsn 天津一女子在路上扶起了摔倒的老人帮忙送去救治并垫付3000元,老人清醒后指认是女子撞到了他,老人的子女要求女子再拿10000元。女子报了警… (A woman in Tianjin helped an elderly person who had fallen on the road, took him for medical treatment, and paid 3,000 yuan. After the elderly person regained consciousness, he accused the woman of causing the accident. The elderly person’s children demanded an additional 10,000 yuan from the woman. The woman reported the incident to the police.) While India is in denial about border tensions in Ladakh, China linked accounts shared this photo with the caption India China border, in which Indian soldiers are seen lying on the ground with hands and legs tied up with rope.
@@
Instruction Now give you a multimodal post and the corresponding textual evidences… image ocr:{CJK}UTF8gbsn 老人跌倒路边一直求救女子暖心扶起帮老人找家我拉你一下好不好@历史课课代表, (The elderly person fell by the roadside, continuously calling for help. The compassionate woman helped the elderly person up and assisted in finding his way home, saying, ”Let me give you a hand, is that okay?” @HistoryClassRep) image caption:an old man lying on the ground with his head down… Evidences… ¡and¿ {CJK}UTF8gbsn 雷音振长空01-0220:57 微博轻享版任何值得称道的正义之光,都应建立在事实之上。可惜的是,这个被这么多大v转发的新闻是假的 (Any commendable light of justice should be built upon facts. Unfortunately, the news widely shared by many influencers is fake.) … Now,you are required to explain… Now give you a multimodal post and the corresponding textual evidences… image ocr: @Lz1277 87829, image caption: soldiers lay on the ground while others stand around them. Evidences… ¡and¿ 72 Indian soldiers injured, captured in clash with Chinese Army… ¡and¿ Footage claims to show ’Indian troops beating a captured Chinese soldier’ amid border tension… ## Explanation: Now, you are required to explain and label the post in three-way classification scheme: [”non-rumor”, ”rumor”, ”unverified”].
@@
gpt-3.5-turbo The post is labeled as rumor because the textual evidences show conflicting information and the image caption does not directly support the text. The textual evidence includes multiple instances of the post being shared on social media, but also includes a comment suggesting that the news is fake. Additionally, the image OCR does not provide any relevant information to confirm the story in the text. Fine-grained labels that are suitable for the post include ”Misleading and False” and ”Fake News” as the story is being disputed and there is a suggestion that it is not true. Therefore, the post is labeled as ¡label¿ rumor ¡/label¿. The post is labeled as a rumor because the textual evidence mentions that the photo was circulated among Chinese military websites, and there are conflicting reports about the incident, with some claiming that Indian soldiers were captured and others claiming that Indian troops beat a captured Chinese soldier. This indicates that there is uncertainty and conflicting information surrounding the photo, leading to its classification as a rumor. As for the fine-grained labels, ”Altered Image” and ”Misleading” would be suitable for the post, as the photo is being shared with a misleading caption and there are conflicting reports about the incident depicted in the image. Therefore, the post is labeled as ¡label¿ rumor ¡/label¿.
@@
$\text{MMIDR}_{\text{LLaMA2}}$ The post is labeled as rumor because the textual and image evidences do not provide clear and credible information about the incident. The textual evidence includes repeated mentions of ”Sina Visitor System,” which is not relevant to the content of the post. The image evidence includes a screenshot of a text message in Chinese, which does not provide any relevant information about the incident. Additionally, the image caption describes an old man lying on the ground, which is not directly related to the claim made in the post. Fine-grained labels that are suitable for the post include ”Misleading,” ”Lacks Context,” and ”Unverified,” as the post lacks clear and credible information, and the image evidence does not support the claim made in the text. Therefore, the post is labeled as ¡label¿ rumor ¡/label¿. The post is labeled as a rumor because the textual evidence suggests that the photo of Indian soldiers being tied up with rope is being circulated among Chinese military websites, but there is no concrete evidence to support this claim. Additionally, the image OCR and caption do not provide any relevant information to verify the claim made in the text. Fine-grained labels that are suitable for the post include: Altered Image, Misleading, Misattributed, and False. Therefore, the post is labeled as ¡label¿ rumor ¡/label¿.
@@
$\text{MMIDR}_{\text{MiniGPT-v2}}$ The post is labeled as a rumor because the textual evidence includes multiple screenshots of Chinese social media posts and text messages, but there is no concrete evidence or credible source to support the claim that the woman in the post helped an old man and paid 3000 yuan for his medical expenses. The image caption of an old man lying on the ground with his head down also does not provide any concrete evidence to support the claim. Fine-grained labels that are suitable for the post include ”Misleading”, ”Lacks Context”, and ”Unverified”. These labels reflect the lack of concrete evidence and the potential for misleading information in the post. Therefore, the post is labeled as ¡label¿ rumor ¡/label¿. The post is labeled as rumor because the textual and image evidences suggest that there is a claim of Indian soldiers being tied up and beaten by Chinese soldiers, which is not verified. The textual evidence mentions a photo circulating among Chinese military websites showing PLA troops bringing down several Indian soldiers at a close-quarters fight at Pangong Lake border area with India. The image evidence does not directly support the claim, but it does not refute it either. Fine-grained labels that are suitable for the post include ”Misleading”, ”Misrepresented”, and ”Unverified”. These labels reflect the lack of verification and the potential for misleading information in the post. Therefore, the post is labeled as ¡label¿ rumor ¡/label¿
@@ 4.4. RQ3: Faithful Detection and Interpretable Results
To evaluate the ability of LLMs in producing high-quality and coherent text explanations for multimodal misinformation, we conducted a case study, as detailed in Table 4.3. We specifically chose one Chinese example and one English example for in-depth analysis.